A. Using the supply and demand equations given below, calculate the market clearing buyers’ and sellers’ prices with no sales tax.

D(P+)= 280 - 2P+ S(P3)= 40 + 4P3
B. Next suppose the government collects a 20% sales tax for each unit sold. Calculate the new market clearing buyers’ and sellers’ price.
Answer #1

N - Demand : a = 280-2P .; Supply : az yot up; p = 0-40 pa 280-6 . A: In equili birum Euantity demanded = Quantity supplied